# Scaling Gaussian Process Regression with Full Derivative Observations

### v1

@@ 4.2 Posterior Inference
1:DSoftKI hyperparameters $\theta=(\ell,\gamma,\mathbf{z},\mathbf{T},\beta_{v},\beta_{g})$.
2:Dataset $(\mathbf{x},\tilde{\mathbf{y}})$.
@@
8:         $\mathbf{F}\leftarrow\tilde{\mathbf{\Sigma}}_{\mathbf{x}_{b}\mathbf{z}}\mathbf{L}$triangleright Low rank representation of $\tilde{\mathbf{K}}_{\mathbf{x}_{b}\mathbf{x}_{b}}^{\text{DSoftKI}}(\ell,\gamma,\mathbf{z},\mathbf{T})=\mathbf{F}\mathbf{F}^{T}$
9:         $\tilde{\mathbf{D}}_{\theta}\leftarrow\texttt{LazyOp}((\texttt{LazyOp}(\mathbf{F}\mathbf{F}^{T})+\texttt{LazyOp}(\tilde{\mathbf{\Lambda}}(\beta_{v},\beta_{g}))))$
10:         $\triangleright$ LazyOp delays computation of result until needed
11:         $\theta\leftarrow\theta+\eta\nabla_{\theta}\log\hat{\tilde{p}}(\tilde{\mathbf{y}}_{b}\,|\,\mathbf{x}_{b};\tilde{\mathbf{D}}_{\theta})$triangleright Stabilized DSoftKI MLL (Equation 22)
@@
13:end for
14:$\alpha\leftarrow\texttt{solve}(\hat{\tilde{\mathbf{C}}}\alpha=\hat{\tilde{\mathbf{K}}}_{\mathbf{z}\mathbf{x}}\tilde{\mathbf{\Lambda}}^{-1}\tilde{\mathbf{y}})$triangleright Solve system of linear equations (Appendix A.2)
15:return $\alpha$
Algorithm 1 DSoftKI Regression. It is the SoftKI algorithm with the DSoftKI kernel in place of the SoftKI kernel.
@@
[…]
where $\tilde{\mathbf{D}}_{\theta}=\tilde{\mathbf{K}}_{\mathbf{x}\mathbf{x}}^{\text{DSoftKI}}(\ell,\gamma,\mathbf{z},\mathbf{T})+\tilde{\mathbf{\Lambda}}(\beta_{v},\beta_{g})$ for $\theta=(\ell,\gamma,\mathbf{z},\mathbf{T},\beta_{v},\beta_{g})$. For a minibatch of size $b$, the resulting matrix is of size $b(d+1)\times b(d+1)$ as opposed to $b\times b$ in SoftKI. For sizable $b$ and $d$, this would be intractable to compute with since it requires solving a system of linear equations.
@@
$\tilde{\mathbf{D}}_{\theta}=\texttt{LazyOp}((\texttt{LazyOp}(\mathbf{F}\mathbf{F}^{T})+\texttt{LazyOp}(\tilde{\mathbf{\Lambda}}(\beta_{v},\beta_{g}))))$ (21)
that directly stores $\mathbf{F}$ and $\tilde{\mathbf{\Lambda}}(\beta_{v},\beta_{g})$, and delays computation of any intermediate result until needed. To take advantage of this representation, we use a low-rank multivariate Gaussian distribution which can avoid working with the full $b(d+1)\times b(d+1)$ matrix $\tilde{\mathbf{D}}_{\theta}$ via the Woodbury matrix identity and matrix determinant lemma. This reduces the challenge of computing the MLL to computing the determinant and inverses of the $m\times m$ capacitance matrix $\mathbf{I}+\mathbf{F}^{T}\tilde{\mathbf{\Lambda}}(\beta_{v},\beta_{g})^{-1}\mathbf{F}$ instead which takes time complexity $\mathcal{O}(m^{2}bd)$ to form. When numerical instability is encountered during the computation of DSoftKI’s MLL, we compute Hutchinson’s pseudoloss. The time complexity of a single MVM is $\mathcal{O}(mbd)$. The stabilizied DSoftKI MLL is thus
@@ D SoftKI SVGP DSVGP DDSVGP DSoftKI
Branin 2 0.004 $\pm$ 0.0 0.016 $\pm$ 0.003 0.081 $\pm$ 0.023 0.183 $\pm$ 0.025 0.003 $\pm$ 0.001
Six-hump-camel 2 0.026 $\pm$ 0.003 0.05 $\pm$ 0.001 0.183 $\pm$ 0.071 0.688 $\pm$ 0.041 0.014 $\pm$ 0.005
Styblinski-tang 2 0.026 $\pm$ 0.001 0.046 $\pm$ 0.006 0.089 $\pm$ 0.023 0.123 $\pm$ 0.043 0.013 $\pm$ 0.004
Hartmann 6 0.05 $\pm$ 0.002 0.163 $\pm$ 0.008 0.335 $\pm$ 0.008 0.338 $\pm$ 0.013 0.01 $\pm$ 0.003
Welch 20 0.01 $\pm$ 0.001 0.065 $\pm$ 0.001 - 0.571 $\pm$ 0.015 0.003 $\pm$ 0.002
Table 2: Test RMSE (best bolded) on selected synthetic datasets .
@@
Branin -0.461 $\pm$ 0.108 0.116 $\pm$ 0.0 -1.463 $\pm$ 0.512 -0.513 $\pm$ 0.166 -4.092 $\pm$ 0.087
Six-hump-camel -0.49 $\pm$ 0.052 0.123 $\pm$ 0.0 -1.535 $\pm$ 0.177 -0.638 $\pm$ 0.05 -2.175 $\pm$ 0.221
Styblinski-tang -0.315 $\pm$ 0.083 0.122 $\pm$ 0.001 -1.197 $\pm$ 0.066 -0.901 $\pm$ 0.121 -2.292 $\pm$ 0.062
Hartmann -0.293 $\pm$ 0.134 0.221 $\pm$ 0.003 -1.369 $\pm$ 0.015 -1.29 $\pm$ 0.044 -3.162 $\pm$ 0.306
Welch -0.662 $\pm$ 0.163 0.141 $\pm$ 0.0 - 0.705 $\pm$ 0.051 -3.742 $\pm$ 0.945
Table 3: Test NLL (best bolded) on selected synthetic datasets.
@@
Branin 0.003 $\pm$ 0.0 -4.159 $\pm$ 0.04 -0.001 -0.243
Six-hump-camel 0.165 $\pm$ 0.023 -0.385 $\pm$ 0.141 0.147 1.746
Styblinski-tang 0.025 $\pm$ 0.006 -1.848 $\pm$ 0.153 0.013 0.455
Hartmann 0.336 $\pm$ 0.035 0.334 $\pm$ 0.096 0.327 3.670
Welch 0.024 $\pm$ 0.009 -1.431 $\pm$ 0.537 0.022 3.253
@@ 5 Experiments
In this section, we evaluate DSoftKI on synthetic functions (Section 5.1) to obtain a baseline of comparison and high-dimensional molecular force field (Section 5.2) to test scale.
Baseline GP and GPwD methods. We use the default GPyTorch implementations (Gardner et al., 2018) of SVGP, DSVGP, and DDSVGP with $2$ inducing directions as baseline variational GPs. We use the PLL (Jankowiak et al., 2020) modification to the ELBO for the variational GPs as recommended by (Padidar et al., 2021) to obtain heteroskedastic noise modeling. We also use SoftKI as a baseline GP where we also adopt the DSoftKI interpolation scheme. We do not compare against DSKI since the dimensionality of the datasets are too high. We use the RBF kernel (with scale) using automatic relevance determination (ARD) lengthscales with the exception of DDSVGP which does not support it. Unless otherwise stated, we use $m=512$ inducing points for all methods. We optimize all hyperparameters using the Adam (Kingma & Ba, 2014) optimizer. All GPs and GPwDs use single-precision floating point numbers, except for DSoftKI which uses double-precision occasionally to enhance the stability of Cholesky decomposition as described previously.
@@ 5.1 Regression with and without Derivative Information
Once we control for the effective learning rate, we see that fitting derivatives helps improve DSoftKI’s performance relative to SoftKI, its non-derivative base, as measured by test RMSE performance. We believe that this increase in performance is due to DSoftKI’s modified interpolation scheme which jointly takes into account distances and directional information via learned temperature hyperparameters. In Table 4, we run DSoftKI with the SoftKI interpolation scheme, and see that it performs worse on several datasets. For the variational methods, we observe that SVGP performs the best, followed by DDSVGP which fits two directions, and followed last by DSVGP which fits all derivative information once effective learning rate is controlled for. As a reminder, we use $\Delta_{\text{eff}}$ to control for the amount of data each method sees during a step of hyperparameter optimization and not as a recommendation to obtain the best possible results. We conjecture that the difference in optimization dynamics between DSoftKI and DSVGP/DDSVGP is because the former jointly models values and gradients with each interpolation point whereas the latter approaches introduces separate inducing points for values and gradients/directions.
Another trend that we observe once we control for $\Delta_{\text{eff}}$ is that GPwDs tend to have lower NLLs compared to their non-derivative counterparts. For methods based on SVGP, this is somewhat surprising since the variants that fit derivatives also have higher test RMSE, indicating that the the versions that fit derivatives result in simpler models. On the other hand, for DSoftKI, we observe that the NLL is positively correlated with its test RMSE relative to the performance of SoftKI. This suggests that the additional information is used to increase the complexity of the model when it also improves the fit of the model.
@@ 5.2 High-Dimensional Molecular Force Fields
In this section, we evaluate the ability of DSoftKI to fit gradient information. Towards this end, we use the MD22 dataset (Chmiela et al., 2023), a dataset of molecular energy surfaces where inputs are molecular configurations ($d=168$ to $1110$), outputs are energies, and gradients of the energy surface correspond to the the physically-meaningful quantity of (negative) forces. Consequently, fitting this information can enable us to model the physical dynamics.
@@ 6 Conclusion
First, while DSoftKI is more scalable than existing GPwD methods that fit derivatives, further work can be done to improve its computational complexity. In particular, the method is currently bottlenecked by the computation of the DSoftKI MLL. Improved approximations of the MLL or a different objective could further scale the efficiency of fitting derivative observations. Second, a deeper investigation of the tradeoffs between fitting derivative information or not when they are available is warranted. For some applications, the added difficulty of fitting gradient observations may not warrant improved surface modeling. Third, it would be an interesting direction fo future work to explore the use of more expressive kernels in DSoftKI such as those based deep kernel learning with DSoftKI. Notably, the construction of the DSoftKI kernel does not require taking its first or second order derivatives, lending itself to tractable hyperparameter optimization with more expressive kernels.

### v2

@@ 4.2 Posterior Inference
1:DSoftKI hyperparameters $\theta=(\ell,\gamma,\mathbf{z}^{(m\times d)},\mathbf{T}^{(m\times d)},\beta_{v},\beta_{g})$.
[…]
@@
8:   $\mathbf{F}\leftarrow\tilde{\mathbf{\Sigma}}_{\mathbf{x}_{b}\mathbf{z}}\mathbf{L}$ $\triangleright$ Low rank representation of $\tilde{\mathbf{K}}_{\mathbf{x}_{b}\mathbf{x}_{b}}^{\text{DSoftKI}}(\ell,\gamma,\mathbf{z},\mathbf{T})=\mathbf{F}\mathbf{F}^{T}$
9:   $\tilde{\mathbf{D}}_{\theta}\leftarrow\llbracket\mathbf{F}\mathbf{F}^{T}+\tilde{\mathbf{\Lambda}}(\beta_{v},\beta_{g})\rrbracket_{z}$ $\triangleright$ $\llbracket\cdot\rrbracket_{z}$ delays computation of result until needed
10:   $\theta\leftarrow\theta+\eta\nabla_{\theta}\log\hat{\tilde{p}}(\tilde{\mathbf{y}}_{b}\,|\,\mathbf{x}_{b};\tilde{\mathbf{D}}_{\theta})$ $\triangleright$ Stabilized DSoftKI MLL (Equation 25)
11:  end for
@@
[…]
14:return $\alpha$
Algorithm 1 DSoftKI regression adapts the SoftKI algorithm to handle the additional computational challenges of the DSoftKI kernel and learning local temperature vectors $\mathbf{T}^{(m\times d)}$.
@@
$\log p(\tilde{\mathbf{y}}|\mathbf{x};\theta)=\mathcal{N}(\tilde{\mathbf{y}}\,|\,0,\tilde{\mathbf{D}}_{\theta})\,.$ (23)
where $\tilde{\mathbf{D}}_{\theta}=\tilde{\mathbf{K}}_{\mathbf{x}\mathbf{x}}^{\text{DSoftKI}}(\ell,\gamma,\mathbf{z},\mathbf{T})+\tilde{\mathbf{\Lambda}}(\beta_{v},\beta_{g})$ for $\theta=(\ell,\gamma,\mathbf{z},\mathbf{T},\beta_{v},\beta_{g})$. We refer the reader to Appendix B.1 for details on hyperparameter initialization. For a minibatch of size $b$, the resulting matrix is of size $b(d+1)\times b(d+1)$ as opposed to $b\times b$ in SoftKI. For sizable $b$ and $d$, this would be intractable to compute with since it requires solving a system of linear equations.
@@
$\tilde{\mathbf{D}}_{\theta}=\llbracket\mathbf{F}\mathbf{F}^{T}+\tilde{\mathbf{\Lambda}}(\beta_{v},\beta_{g})\rrbracket_{z}$ (24)
that directly stores $\mathbf{F}$ and $\tilde{\mathbf{\Lambda}}(\beta_{v},\beta_{g})$, and $\llbracket\cdot\rrbracket_{z}$ delays computation of any intermediate result until needed. To take advantage of this representation, we use a low-rank multivariate Gaussian distribution which can avoid working with the full $b(d+1)\times b(d+1)$ matrix $\tilde{\mathbf{D}}_{\theta}$ via the Woodbury matrix identity and matrix determinant lemma. This reduces the challenge of computing the MLL to computing the determinant and inverses of the $m\times m$ capacitance matrix $\mathbf{I}+\mathbf{F}^{T}\tilde{\mathbf{\Lambda}}(\beta_{v},\beta_{g})^{-1}\mathbf{F}$ instead which takes time complexity $\mathcal{O}(m^{2}bd)$ to form. When numerical instability is encountered during the computation of DSoftKI’s MLL, we compute Hutchinson’s pseudoloss. The time complexity of a single MVM is $\mathcal{O}(mbd)$. The stabilized DSoftKI MLL is thus
@@
Appendix B.2 contains an ablation of DSoftKI using the DSoftKI MLL compared to the Hutchinson’s pseudoloss.
@@ 4.3 The Role of Value and Gradient Noises
Branin 2 0.004 $\pm$ 0.0 0.018 $\pm$ 0.004 0.088 $\pm$ 0.062 0.176 $\pm$ 0.031 0.003 $\pm$ 0.001
Six-hump-camel 2 0.026 $\pm$ 0.003 0.05 $\pm$ 0.003 0.101 $\pm$ 0.031 0.669 $\pm$ 0.04 0.015 $\pm$ 0.006
Styblinski-tang 2 0.025 $\pm$ 0.001 0.05 $\pm$ 0.003 0.101 $\pm$ 0.01 0.125 $\pm$ 0.06 0.012 $\pm$ 0.002
Hartmann 6 0.05 $\pm$ 0.002 0.164 $\pm$ 0.006 0.335 $\pm$ 0.009 0.346 $\pm$ 0.019 0.011 $\pm$ 0.004
Welch 20 0.01 $\pm$ nan 0.065 $\pm$ 0.001 - 0.578 $\pm$ 0.013 0.003 $\pm$ 0.003
Table 2: Test RMSE (best bolded) on selected synthetic datasets. One of the runs for SoftKI encountered numerical instability (hence $\pm\texttt{nan}$). - indicates a timeout for DSVGP.
@@
Branin -0.725 $\pm$ 0.017 0.116 $\pm$ 0.0 -1.044 $\pm$ 0.885 -0.532 $\pm$ 0.068 -4.432 $\pm$ 0.162
Six-hump-camel -0.743 $\pm$ 0.057 0.123 $\pm$ 0.001 -1.539 $\pm$ 0.159 -0.64 $\pm$ 0.058 -2.175 $\pm$ 0.237
Styblinski-tang -0.54 $\pm$ 0.082 0.123 $\pm$ 0.001 -1.182 $\pm$ 0.166 -1.064 $\pm$ 0.188 -2.296 $\pm$ 0.045
Hartmann -0.494 $\pm$ 0.111 0.221 $\pm$ 0.001 -1.362 $\pm$ 0.029 -1.324 $\pm$ 0.027 -3.16 $\pm$ 0.325
Welch -0.699 $\pm$ nan 0.141 $\pm$ 0.0 - 0.706 $\pm$ 0.055 -1.14 $\pm$ 4.795
Table 3: Test NLL (best bolded) on selected synthetic datasets. One of the runs for SoftKI encountered numerical instability (hence $\pm\texttt{nan}$). - indicates a timeout for DSVGP.
d SoftKI SVGP DSVGP DDSVGP DSoftKI
Branin 2 0.194 $\pm$ 0.005 0.339 $\pm$ 0.149 1.318 $\pm$ 0.02 1.945 $\pm$ 0.022 2.555 $\pm$ 0.199
Six-hump-camel 2 0.257 $\pm$ 0.092 0.373 $\pm$ 0.104 1.31 $\pm$ 0.008 1.955 $\pm$ 0.062 2.434 $\pm$ 0.087
Styblinski-tang 2 0.21 $\pm$ 0.006 0.335 $\pm$ 0.14 1.316 $\pm$ 0.018 1.91 $\pm$ 0.016 2.539 $\pm$ 0.137
Hartmann 6 0.199 $\pm$ 0.003 0.337 $\pm$ 0.126 9.305 $\pm$ 0.113 1.934 $\pm$ 0.045 2.544 $\pm$ 0.091
Welch 20 0.213 $\pm$ 0.002 0.418 $\pm$ 0.123 - 1.944 $\pm$ 0.063 4.452 $\pm$ 0.192
Table 4: Wall-clock training time in seconds per epoch (best GPwD bolded) on selected synthetic dataset. - indicates a timeout for DSVGP.
@@
Branin 0.002 $\pm$ 0.0 -4.266 $\pm$ 0.009 -0.000 0.166
Six-hump-camel 0.163 $\pm$ 0.02 -0.392 $\pm$ 0.112 0.148 1.783
Styblinski-tang 0.026 $\pm$ 0.002 -1.829 $\pm$ 0.171 0.014 0.467
Hartmann 0.336 $\pm$ 0.034 0.336 $\pm$ 0.092 0.326 3.495
Welch 0.027 $\pm$ 0.008 -1.311 $\pm$ 0.405 0.024 -0.171
@@ 5 Experiments
In this section, we evaluate DSoftKI on synthetic functions (Section 5.1) to obtain a baseline of comparison and high-dimensional molecular force field (Section 5.2) to test scale. We refer the reader to Appendix C for more experiments with a toy n-body simulation and on the UCI dataset (Dua and Graff, 2017) with synthetic gradients.
Baseline GP and GPwD methods. We use the default GPyTorch implementations (Gardner et al., 2018) of SVGP, DSVGP, and DDSVGP with $2$ inducing directions as baseline variational GPs. We use the PLL (Jankowiak et al., 2020) modification to the ELBO for the variational GPs as recommended by (Padidar et al., 2021) to allow separate noise parameters for function values ($\beta_{v}^{2}$) and gradients ($\beta_{g}^{2}$). We also use SoftKI as a baseline GP where we also adopt the DSoftKI interpolation scheme. We do not compare against DSKI since the dimensionality of the datasets are too high. We use the RBF kernel (with scale) using automatic relevance determination (ARD) lengthscales with the exception of DDSVGP which does not support it. Unless otherwise stated, we use $m=512$ inducing points for all methods. We optimize all hyperparameters using the Adam (Kingma and Ba, 2014) optimizer. All GPs and GPwDs use single-precision floating point numbers, except for DSoftKI which uses double-precision occasionally to enhance the stability of Cholesky decomposition as described previously.
@@ 5.1 Regression with and without Derivative Information
n d SoftKI SVGP DDSVGP DSoftKI
Ac-ala3-nhme 76598 126 0.894 $\pm$ 0.006 1.051 $\pm$ 0.018 7.907 $\pm$ 0.004 82.665 $\pm$ 0.06
Dha 62777 168 0.887 $\pm$ 0.021 0.951 $\pm$ 0.027 6.544 $\pm$ 0.012 89.155 $\pm$ 0.223
At-at 18000 180 0.336 $\pm$ 0.006 0.427 $\pm$ 0.001 2.626 $\pm$ 0.01 27.236 $\pm$ 0.037
Stachyose 24544 261 0.475 $\pm$ 0.005 0.555 $\pm$ 0.007 3.344 $\pm$ 0.021 54.181 $\pm$ 0.142
At-at-cg-cg 9137 354 0.355 $\pm$ 0.008 0.423 $\pm$ 0.007 1.988 $\pm$ 0.006 27.102 $\pm$ 0.036
Buckyball-catcher 5491 444 0.244 $\pm$ 0.007 0.295 $\pm$ 0.004 1.249 $\pm$ 0.007 20.764 $\pm$ 0.111
Double-walled 4528 1110 0.409 $\pm$ 0.007 0.434 $\pm$ 0.004 1.906 $\pm$ 0.003 45.379 $\pm$ 0.007
Table 6: Wall-clock training time in seconds per epoch (best GPwD bolded) on MD22 dataset. We abbreviate Double-walled-nanotube (Figure 4) as Double-walled due to space.
22footnotetext: dfdf
@@
Once we control for the effective learning rate, we see that fitting derivatives helps improve DSoftKI’s performance relative to SoftKI, its non-derivative base, as measured by test RMSE performance. We believe that this increase in performance is due to DSoftKI’s modified interpolation scheme which jointly takes into account distances and directional information via learned temperature hyperparameters. In Table 5, we run DSoftKI with the SoftKI interpolation scheme, and see that it performs worse on several datasets. For the variational methods, we observe that SVGP performs the best, followed by DDSVGP which fits two directions, and followed last by DSVGP which fits all derivative information once effective learning rate is controlled for. As a reminder, we use $\Delta_{\text{eff}}$ to control for the amount of data each method sees during a step of hyperparameter optimization and not as a recommendation to obtain the best possible results. We conjecture that the difference in optimization dynamics between DSoftKI and DSVGP/DDSVGP is because the former jointly models values and gradients with each interpolation point whereas the latter approaches introduce separate inducing points for values and gradients/directions.
Another trend that we observe once we control for $\Delta_{\text{eff}}$ is that GPwDs tend to have lower NLLs compared to their non-derivative counterparts. For methods based on SVGP, this is somewhat surprising since the variants that fit derivatives also have higher test RMSE, indicating that the versions that fit derivatives result in simpler models. On the other hand, for DSoftKI, we observe that the NLL is positively correlated with its test RMSE relative to the performance of SoftKI. This suggests that the additional information is used to increase the complexity of the model when it also improves the fit of the model.
Table 4 reports the (wall-clock) training time (in seconds) per epoch of training on selected synthetic datasets. One epoch of training is one iteration through the dataset. We use a single RTX 6000 Blackwell Pro Max-Q edition GPU. We observe that the non-derivative methods such as SoftKI and SVGP are faster than the derivative-fitting counterparts. This is expected since we fit $d$ times more data in GPwD regression. For methods that predict full derivative observations, we see that DSoftKI has higher per-epoch training time compared to DSVGP, another method that can predict full derivative observations, on datasets with $d=2$. However, as $d$ scales, DSVGP’s cubic scaling in $d$ leads to prohibitive cost and it becomes infeasible to run DSVGP on the Welch ($d=20$) dataset. DSoftKI also has higher per-epoch training time than DDSVGP, particularly for large d. This is expected given DSoftKI’s $O(m^{2}nd)$ complexity versus DDSVGP’s $O(m^{3}p^{3})$ with $p=2$ so that its per-epoch time does not scale with increasing $d$. However, DDSVGP only predicts $p=2$ directional derivatives.
@@ 5.2 High-Dimensional Molecular Force Fields
In this section, we evaluate the ability of DSoftKI to fit gradient information. Towards this end, we use the MD22 dataset (Chmiela et al., 2023), a dataset of molecular energy surfaces where inputs are molecular configurations ($d=168$ to $1110$), outputs are energies, and gradients of the energy surface correspond to the physically-meaningful quantity of (negative) forces. Consequently, fitting this information can enable us to model the physical dynamics.
@@
Table 6 reports the (wall-clock) training time (in seconds) per epoch of training on the MD22 dataset. We include SoftKI and SVGP as non-derivative methods for completeness. As a reminder, DVSGP does not scale to this setting and we use DDSVGP with $p=2$ inducing directions. As expected, DSoftKI has higher per-epoch training time than DDSVGP, particularly for large d. The additional time complexity is justified in cases such as molecular dynamics where full forces are required since the shape of the surface also needs to be fit. Notably, DSoftKI’s training time scales roughly linearly with $d$ (e.g., $20.8$ seconds at $d=444$ versus $45.4$ seconds at $d=1110$), consistent with its $O(m^{2}nd)$ complexity. In contrast, DDSVGP’s time remains nearly constant across dimensions since its $O(m^{3}p^{3})$ complexity is independent of $d$. However, this assumes $mp^{2}>d$, which we observe begins to break when comparing its timing on the Double-walled dataset with the Buckyball-catcher dataset.
@@ 6 Conclusion
First, while DSoftKI is more scalable than existing GPwD methods that fit derivatives, further work can be done to improve its computational complexity. In particular, the method is currently bottlenecked by the computation of the DSoftKI MLL. Improved approximations of the MLL or a different objective could further scale the efficiency of fitting derivative observations. Second, a deeper investigation of the tradeoffs between fitting derivative information or not when they are available is warranted. For some applications, the added difficulty of fitting gradient observations may not warrant improved surface modeling. Third, while we demonstrate that DSoftKI supports DKL out of the box (Appendix D), achieving computational speedups by projecting into a lower-dimensional feature space, further exploration of learned kernels for GPwD regression is a promising direction.